What is the potential energy of 800-kg elevator at the top of a 380 meter tall building?


1
Expert's answer
2021-04-15T07:18:24-0400

The potential energy is given by.


"U_{g}=m\;\;g\;h"


Where.

"h=380\;\text{m}" is the height above the ground.

"g=9.81\;\text{m}/\text{s}^{2}" is the acceleration of gravity.

"m=800\;\text{Kg}" is the mass


Evaluating numerically.


"U_{g}=m\;\;g\;h\\
U_{g}=800\;\text{Kg}\times 9.81\;\text{m}/\text{s}^{2}\times 380\;\text{m}\\
U_{g}=2,982,240\;\text{J}
\\"


Expressing in scientific notation with three significant figures.


"U_{g}=2.98\times 10^{6}\;\text{J}"


The potential energy is "\displaystyle \color{red}{\boxed{U_{g}=2.98\times 10^{6}\;\text{J}}}"